# Experimentally distinguishable origin for electroweak symmetry breaking

### v2

@@ III Numerical Analysis of the Higgs boson trilinear coupling
In our conformal model, once a value for $M_{h_{2}}$ is fixed, we can determine the $U(1)_{H}$ gauge coupling $g_{H}$ and $Z^{\prime}$ boson mass $m_{Z^{\prime}}$ as functions of $|\theta|$. In Fig. 3, we show the relation between $g_{H}$, $m_{Z^{\prime}}$, and $\sin^{2}(\theta)$ for $M_{h_{2}}=10$ GeV, 25 GeV, and 50 GeV, from bottom to top. If the $Z^{\prime}$ boson couples to SM fermions, we may consider high-energy collider bounds on the $(m_{Z^{\prime}},g_{H})$ parameter space. As a well known example of such a model, one may consider the minimal $B-L$ (baryon number minus lepton number) model, based on the $U(1)_{B-L}$ gauge group Davidson:1978pm ; Mohapatra:1980qe ; Marshak:1979fm ; Wetterich:1981bx ; Masiero:1982fi ; Mohapatra:1982xz ; Buchmuller:1991ce . The final results from the LHC Run 2 for the $B-L$ $Z^{\prime}$ gauge boson resonance search provide us with an upper bound on $g_{H}=g_{B-L}$ as a function of $m_{Z^{\prime}}$, which is presented in Ref. Das:2019fee . In Fig. 4, we show the relationship between $g_{H}$ and $m_{Z^{\prime}}$, along with the excluded region (grey shaded) by the Large Hadron Collider (LHC) experiments from Ref. Das:2019fee , if we identify our $U(1)_{H}$ with $U(1)_{B-L}$. The three lines show the relationship for $M_{h_{2}}=$ 10 GeV (red), 25 GeV (black), and 50 GeV (blue) for the range of $0.001\leq\sin^{2}(\theta)\leq 0.8$, corresponding to the range shown in Fig. 2. The lines will extend to the right as we take lower values of $\sin^{2}(\theta)<0.001$. To be consistent with the LHC bounds, $\sin^{2}(\theta)$ must be much smaller than $0.001$ (see Fig. 3). Thus, to explore interesting regions for Higgs phenomenology ($\sin^{2}(\theta)\gtrsim 0.002$), we do not consider the $B-L$ model identification of $U(1)_{H}$. However, note that the identification of $U(1)_{H}$ with well-known flavor-dependent $U(1)$ extended SMs, such as $U(1)_{B-L_{3}}$ Babu:2017olk ; Alonso:2017uky ; Bian:2017rpg and $U(1)_{L_{\mu}-L_{\tau}}$ He:1990pn ; Foot:1990mn , is still possible, as in these models the $Z^{\prime}$ boson has no coupling with the first generation quarks and can easily evade the LHC constraints.
@@ IV Conclusion and discussion
We have considered a classically conformal $U(1)$ extension of the SM where the $U(1)$ Higgs field ($\Phi$) with a $U(1)$ charge $+2$ is introduced. The $U(1)$ symmetry is radiatively broken by the Coleman-Weinberg mechanism, which generates a negative mass squared to the SM Higgs doublet through the mixed quartic coupling in the scalar potential $V\supset-\lambda_{mix}(\Phi^{\dagger}\Phi)(H^{\dagger}H)$, and as a result, the EW symmetry is broken. Therefore, the radiative symmetry breaking of the new U(1) gauge symmetry is the origin of the EW symmetry breaking. This is a crucial difference from the conventional Higgs potential in which the EW symmetry breaking is triggered by a negative mass squared introduced by hand at the tree-level. In order to investigate the Higgs phenomenology of the classically conformal model, we have analyzed the effective Higgs potential and read off the trilinear coupling of the SM-like Higgs boson ($h_{1}$) with a pair of the SM singlet-like Higgs bosons ($h_{2}$) for the case of $M_{h_{1}}>2M_{h_{2}}$. For comparison, we have also calculated the trilinear coupling for the conventional Higgs potential. What we have found is very intriguing: for the same mixing angle between $H$ and $\Phi$ and the same Higgs boson mass spectrum, the trilinear coupling in the conformal model is highly suppressed compared to the one from the conventional Higgs potential, which is likely a striking nature of the classically conformal potential. We then have considered experimental signals for such conformal structure via anomalous SM Higgs boson couplings from the mixing between two Higgs fields and the anomalous (SM-like) Higgs boson decay $h_{1}\rightarrow h_{2}h_{2}$ followed by $h_{2}\to b\bar{b}$. Future $e^{+}e^{-}$ colliders, in particular, the proposed ILC will operate as a Higgs factory which allows us to precisely measure the SM Higgs boson properties. The result of our analysis indicates that once the anomalous SM-like Higgs boson coupling is measured at the ILC, the observation of anomalous Higgs boson decay $h_{1}\to h_{2}h_{2}$ is promising in the conventional Higgs potential, while this decay process is highly suppressed and not detectable in the classically conformal model. This is a way to to distinguish the origin of EW symmetry breaking.

### v3

@@ III Numerical Analysis of the Higgs boson trilinear coupling
In our conformal model, once a value for $M_{h_{2}}$ is fixed, we can determine the $U(1)_{H}$ gauge coupling $g_{H}$ and $Z^{\prime}$ boson mass $m_{Z^{\prime}}$ as functions of $|\theta|$. In Fig. 3, we show the relation between $g_{H}$, $m_{Z^{\prime}}$, and $\sin^{2}(\theta)$ for $M_{h_{2}}=10$ GeV, 25 GeV, and 50 GeV, from bottom to top. If the $Z^{\prime}$ boson couples to SM fermions, we may consider high-energy collider bounds on the $(m_{Z^{\prime}},g_{H})$ parameter space. As a well known example of such a model, one may consider the minimal BLB-L (baryon number minus lepton number) model, based on the $U(1)_{B-L}$ gauge group Davidson:1978pm ; Mohapatra:1980qe ; Marshak:1979fm ; Wetterich:1981bx ; Masiero:1982fi ; Mohapatra:1982xz ; Buchmuller:1991ce . The final results from the LHC Run 2 for the BLB-L $Z^{\prime}$ gauge boson resonance search provide us with an upper bound on $g_{H}=g_{B-L}$ as a function of $m_{Z^{\prime}}$, which is presented in Ref. Das:2019fee . In Fig. 4, we show the relationship between $g_{H}$ and $m_{Z^{\prime}}$, along with the excluded region (grey shaded) by the Large Hadron Collider (LHC) experiments from Ref. Das:2019fee , if we identify our $U(1)_{H}$ with $U(1)_{B-L}$. The three lines show the relationship for $M_{h_{2}}=$ 10 GeV (red), 25 GeV (black), and 50 GeV (blue) for the range of $0.001\leq\sin^{2}(\theta)\leq 0.8$, corresponding to the range shown in Fig. 2. The lines will extend to the right as we take lower values of $\sin^{2}(\theta)<0.001$. To be consistent with the LHC bounds, $\sin^{2}(\theta)$ must be much smaller than $0.001$ (see Fig. 3). Thus, to explore interesting regions for Higgs phenomenology ($\sin^{2}(\theta)\gtrsim 0.002$), we do not consider the BLB-L model identification of $U(1)_{H}$. However, note that the identification of $U(1)_{H}$ with well-known flavor-dependent $U(1)$ extended SMs, such as $U(1)_{B-L_{3}}$ Babu:2017olk ; Alonso:2017uky ; Bian:2017rpg and $U(1)_{L_{\mu}-L_{\tau}}$ He:1990pn ; Foot:1990mn , is still possible, as in these models the $Z^{\prime}$ boson has no coupling with the first generation quarks and can easily evade the LHC constraints.
@@ IV Conclusion and discussion
We have considered a classically conformal $U(1)$ extension of the SM where the $U(1)$ Higgs field ($\Phi$) with a $U(1)$ charge +2+2 is introduced. The $U(1)$ symmetry is radiatively broken by the Coleman-Weinberg mechanism, which generates a negative mass squared to the SM Higgs doublet through the mixed quartic coupling in the scalar potential $V\supset-\lambda_{mix}(\Phi^{\dagger}\Phi)(H^{\dagger}H)$, and as a result, the EW symmetry is broken. Therefore, the radiative symmetry breaking of the new U(1) gauge symmetry is the origin of the EW symmetry breaking. This is a crucial difference from the conventional Higgs potential in which the EW symmetry breaking is triggered by a negative mass squared introduced by hand at the tree-level. In order to investigate the Higgs phenomenology of the classically conformal model, we have analyzed the effective Higgs potential and read off the trilinear coupling of the SM-like Higgs boson ($h_{1}$) with a pair of the SM singlet-like Higgs bosons ($h_{2}$) for the case of $M_{h_{1}}>2M_{h_{2}}$. For comparison, we have also calculated the trilinear coupling for the conventional Higgs potential. What we have found is very intriguing: for the same mixing angle between $H$ and $\Phi$ and the same Higgs boson mass spectrum, the trilinear coupling in the conformal model is highly suppressed compared to the one from the conventional Higgs potential, which is likely a striking nature of the classically conformal potential. We then have considered experimental signals for such conformal structure via anomalous SM Higgs boson couplings from the mixing between two Higgs fields and the anomalous (SM-like) Higgs boson decay $h_{1}\rightarrow h_{2}h_{2}$ followed by $h_{2}\to b\bar{b}$. Future $e^{+}e^{-}$ colliders, in particular, the proposed ILC will operate as a Higgs factory which allows us to precisely measure the SM Higgs boson properties. The result of our analysis indicates that once the anomalous SM-like Higgs boson coupling is measured at the ILC, the observation of anomalous Higgs boson decay $h_{1}\to h_{2}h_{2}$ is promising in the conventional Higgs potential, while this decay process is highly suppressed and not detectable in the classically conformal model. This is a way to to distinguish the origin of EW symmetry breaking.